# Chirality-induced selectivity of angular momentum by orbital Edelstein effect in carbon nanotubes

### v1

@@ Abstract
CNTs can be constructed mathematically by ‘cutting’ a rectangle out of the two-dimensional carbon allotrope graphene and rolling it to a cylinder. The rotation angle $\alpha$ between the rectangle and the underlying graphene lattice influences the resulting physical properties; most famously, some tubes are metallic and others are insulting. While the two most well-known types of CNTs – tubes with armchair or zigzag edges – posses inversion symmetry, all other nanotubes characterized by $0^{\circ}<|\alpha|<30^{\circ}$ have a broken inversion symmetry which makes them chiral. Over the recent years, such chirality-pure nanotubes have become more easily available (review see Ref. [4]), such as the chiral $(12,6)$ CNTs that were grown with a Co7W6 or Mo2C catalyst with a purity of more than $90\%$ [13, 14]. Furthermore, dispersion by surfactants [15] and other methods [4] allow for sorting grown tubes based on their chirality.
Figure 1: Chiral carbon nanotube. a Geometrical construction of a $(n,m)=(7,1)$ nanotube from a two-dimensional graphene lattice. The color indicates the $x$ coordinate of each atom. The circumferential vector $\boldsymbol{C}$ and the translational vector $\boldsymbol{T}$ span a rectangle containing the atoms of the unit cells of the corresponding chiral carbon nanotube in b. The chirality angle $\alpha$ is indicated. c Top view of the nanotube.
@@
Figure 2: Chirality-induced orbital selectivity of a chiral carbon nanotube. a Band structure of a metallic $(n,m)=(7,1)$ nanotube for which the color indicates the value of the orbital angular momentum $L_{\nu,z}(k)$ normalized to a unit tube length (see legend). b Orbital Edelstein susceptibility $\chi_{z}^{L_{z}}\cdot a/T$ as a function of energy for a $(n,m)=(7,1)$ nanotube (blue) and for the corresponding nanotube with opposite chirality $(n,m)=(8,-1)$ (red). The dashed line indicates an energy of $E=0.49\,\mathrm{eV}$ in the range where the orbital Edelstein susceptibility increases quadratically with energy. c Band structure and d orbital Edelstein susceptibility (blue) of an insulating $(n,m)=(4,2)$ nanotube and for the corresponding nanotube with opposite chirality $(n,m)=(6,-2)$ (red).
Over the past few years, the orbital Hall effect [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] and the orbital Edelstein effect [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56] have been established as the counterparts to the spin Hall effect [57, 58] and spin Edelstein effect [59, 60]. It has been calculated that the orbital Edelstein effect might explain the large measured CISS: It might mostly be caused by chirality-induced orbital selectivity (CIOS) that is only converted to CISS upon measurement [47, 48, 53, 54, 56, 61]. The orbital angular momentum is transformed to spin angular momentum due to the spin-orbit coupling in the electrodes. In a recent study, we have shown that CIOS in a helix structure like in tellurium has a purely geometrical origin and can be explained even classically based on a helical trajectory of a moving electron [61]. Orbital angular momentum is generated by a rotational motion and is transported along the helix. The effect is caused by intersite-hybridization of the wave function and has to be calculated using the modern formulation [62, 63, 64, 65, 66] instead of relying on the often-used atomic-center approximation. So far, the effect has not been analyzed for more complicated structures, like CNTs, where the sub-band structure and Dirac physics make a classical comparison questionable.
@@ Results
Here we consider ideal CNTs that are periodic along the $z$ direction. The unit cell can be constructed geometrically by rolling up a rectangular cut of a two-dimensional graphene layer. This layer is characterized by two basis atoms and the two lattice vectors $\boldsymbol{a}_{1}=a\boldsymbol{e}_{x}$ and $\boldsymbol{a}_{2}=(a/2)\boldsymbol{e}_{x}-(a\sqrt{3}/2)\boldsymbol{e}_{y}$ with the lattice constant $a=2.46\,$Å.
@@
The rectangle that ultimately forms the unit cell of the nanotube is spanned by the circumferential vector $\boldsymbol{C}=n\boldsymbol{a}_{1}+m\boldsymbol{a}_{2}$ and the translational vector $\boldsymbol{T}=\frac{2m+n}{d}\boldsymbol{a}_{1}-\frac{2n+m}{d}\boldsymbol{a}_{2}$ with $d$ the greatest common divisor of $2m+n$ and $2n+m$. The values of $n\in\mathbb{N}^{+}$ and $m=0,1,\dots,n$ determine the configuration of the tube. If the tube was cut horizontally, a $(n,m)=(n,0)$ tube would have a zigzag edge and a $(n,m)=(n,n)$ tube would have an armchair edge. Both of these types of tubes are achiral but all other tubes, for which $m\neq 0$ and $m\neq n$, are chiral. Note that tubes with the opposite chirality can be constructed by allowing also for negative values of $m$.
@@
Figure 3: Band structure of a metallic carbon nanotube by backfolding the graphene band structure. a Brillouin zone of two-dimensional graphene (hexagon) and smaller Brillouin zone (gray) corresponding to the unit cell spanned by $\boldsymbol{C}$ and $\boldsymbol{T}$ (gray rectangle in Fig. 1a) for a $(n,m)=(7,1)$ tube. Black lines represent multiples of the reciprocal lattice vector $\boldsymbol{b}_{T}$ shifted by multiples of the reciprocal lattice vector $\boldsymbol{b}_{C}$. b Band structure of two-dimensional graphene. The black lines from panel a give rise to the band structure of the corresponding nanotube in c once they are shifted back to the small Brillouin zone (gray). The color resembles $L_{\nu,z}^{\mathrm{approx}}$ that approximates $L_{\nu,z}(k)$ of the corresponding carbon nanotube. The quantity has been normalized by $a/T$ for comparability. Colored dots indicate the Dirac points originally located at the points $K$ and $K^{\prime}$ of the Brillouin zone of graphene.
@@
An example of the geometrical construction of a chiral CNT is shown in Fig. 1. The circumferential vector shown in 1a is $\boldsymbol{C}=7\frac{1}{2}a\boldsymbol{e}_{x}-(a\sqrt{3}/2)\boldsymbol{e}_{y}=7\boldsymbol{a}_{1}+1\boldsymbol{a}_{2}$. The gray rectangle is rolled up to form the unit cell of the $(7,1)$ CNT in Fig. 1b. This nanotube is chiral with a chirality angle of
@@
$\alpha=\sphericalangle(\boldsymbol{T}_{\mathrm{graphene}},\boldsymbol{e}_{y})=\arctan\frac{-m\sqrt{3}/2}{n+m/2},$ (1)
@@
in our example $\alpha\approx-6.59^{\circ}$. The angle can be recognized by following the color code in Fig. 1b along the translational direction $\boldsymbol{T}$. When we look into the tube from the top (cf. Fig. 1c), this color code results in a spiral as a signature of the chirality. Due to the 6-fold rotational symmetry of graphene, we will analyze only configurations with $|\alpha|\leq 30^{\circ}$ and focus on the ones with negative chirality angles. Note, that the definition of the chirality angle $\alpha$ varies throughout the literature and can be defined with the opposite sign by choosing a different reference line of the graphene lattice.
@@
Diagonalizing the tight-binding Hamiltonian results in the band structure and eigenvectors. The band structure is symmetric and exhibits many bands $E_{\nu}(k)$ due to the large unit cell. Since the tube is only periodic along the tube vector $\boldsymbol{T}\parallel\boldsymbol{e}_{z}$, we find dispersion only along $k\equiv k_{z}$. The system behaves quasi-one-dimensional. Note, however, that the chirality of the system is accounted for by the hopping paths along the $x$ and $y$ directions.
@@
The band structure of our example nanotube $(n,m)=(7,1)$ (cf. Fig. 2a) consists of 76 bands and can be easily understood based on backfolding the graphene band structure (cf. Fig. 3). Most importantly, the two Dirac points at the Fermi level of graphene exist also for this nanotube. This is the case for all metallic nanotubes, $\mathrm{mod}(n-m,3)=0$. All other nanotubes exhibit a band gap and have no Dirac points near the Fermi level and are characterized by $\mathrm{mod}(n-m,3)=1,2$; an example is shown in Fig. 2c.
@@
In Fig. 2(a,c), the orbital angular momentum is added as a colorcode (normalized to a comparable length, $L_{\nu,z}(k)\cdot a/T$). It has been calculated based on the modern formulation of orbital magnetization (for details see Methods). For the considered metallic nanotube, the values are in the order of up to ±$\sim\pm 0.5\,\hbar$ and depend antisymmetrically on $k$. $L_{\nu,z}$ is exactly zero at the energy of the Dirac point. In fact, if we expand the system linearly around this energy, the orbital angular momentum is zero everywhere $L_{\nu,z}(k)\equiv 0$. This can be interpreted as a consequence of the massless Dirac fermions. However, the actual band structure deviates from a perfect Dirac cone. As a consequence, $L_{\nu,z}(k)$ increases approximately linearly with $(k-k_{\mathrm{Dirac}})$. However, this dependence is slightly different for $k>k_{\mathrm{Dirac}}$ and $k<k_{\mathrm{Dirac}}$. For the other Dirac point, the orbital angular momentum is exactly reversed. This delicate $k$-dependence determines the orbital Edelstein effect in this energy range, as we will discuss later.
@@
The orbital angular momentum is caused by the circular motion in the nanotube. Coming back to the two-dimensional graphene layer that forms the nanotube, this motion corresponds to a translational motion along the direction $\boldsymbol{C}$. Making use of the semiclassical formula $\boldsymbol{L}=\boldsymbol{r}\times\boldsymbol{p}$, we find for each band of the nanotube that $L_{\nu,z}$ is approximately given by the product of the radius of the tube $r=\nicefrac{{C}}{{2\pi}}$, the electron mass $m_{e}$ and the velocity $v_{C}=\frac{1}{\hbar}\frac{\partial E_{\mathrm{2d}}(k_{T},k_{C})}{\partial k_{C}}$ along the $\boldsymbol{C}$ direction which is calculated from the band structure $E_{\mathrm{2d}}(k_{T},k_{C})$ of the graphene layer as $L_{\nu,z}^{\mathrm{approx}}(k)\sim m_{e}rv_{C}(k_{T},k_{C})=\frac{m_{e}}{\hbar}\frac{C}{2\pi}\frac{\partial E_{\mathrm{2d}}(k_{T},k_{C})}{\partial k_{C}}$. Note that $k_{T}$ corresponds to $k$ in the nanotube and $k_{C}$ determines the sub-band index $\nu$.
@@
The orbital Edelstein susceptibility per unit length of the tube $\chi_{z}^{L_{z}}\cdot a/T$ (Fig. 2b) is considerably large in that energy range with values up to $\sim 10^{-8}\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. At $E=0.49\,\mathrm{eV}$, (dashed line in Fig. 2) $\chi_{z}^{L_{z}}\cdot a/T\sim 10^{-9}\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. For energies farther away from the location of the Fermi level $E_{\mathrm{F}}=0$, we find oscillations with energy. Here, the orbital Edelstein susceptibility even reaches values of $\sim 10^{-7}\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. The oscillations can be understood based on the formation of the sub-band structure by backfolding, shown in Fig.  3. The backfolding leads to different oscillations of the orbital Edelstein susceptibility for different metallic CNTs except for the energy region close to the Fermi level. Here, there are always only four bands that resemble the two Dirac points. The orbital Edelstein susceptibility for different metallic CNTs, including the (12,6) CNTs mentioned in the introduction, is shown in Fig. S1a in the Supplementary Information (SI) and will be systematically analyzed below.
Figure 4: Comparison of metallic carbon nanotubes. The diagram shows all metallic nanotubes that are characterized by a circumference $C\leq 14a$ (plotted along the radial direction) and a chirality angle $|\alpha|\leq 30^{\circ}$ (plotted along the polar direction). The color of the hexagon indicates the value of the orbital Edelstein susceptibility at an energy of $E=0.49\,\mathrm{eV}$ (dashed line in Fig. 2). This value has been normalized by the length of the tube $T$, for comparability: $\chi_{z}^{L_{z}}\cdot a/T$ (see legend). The color in the background is the function proportional to $C\sin(\alpha/6)$.
Here we want to note that for the insulating CNTs, the backfolded band structure, shown in Fig. S2 in the Supplementary Information (SI), reveals that there are no Dirac points and that there is a band gap at the Fermi level. Still, the origin of the finite orbital Edelstein susceptibility is the same as for the metallic cases: We see a finite orbital angular momentum that is antisymmetric with $k$, as presented in Fig. 2c for a $(n,m)=(4,2)$ tube. Like for the metallic cases, there are oscillations in the energy-dependent orbital Edelstein susceptibility (Fig. 2d) that can be explained based on the backfolding. A comparison of various other insulating CNTs is shown in Fig. S1b in the Supplementary Information (SI).
@@
Figure 5: Chirality dependence of the orbital Edelstein susceptibility per atom. a The blue points are the calculated values $\chi_{z,\mathrm{atom}}^{L_{z}}$ at an energy of $E=0.49\,\mathrm{eV}$ for all metallic nanotubes shown in Fig. 4. Here the band structure exhibits four nearly linear bands and the effect is governed by the chirality of the tube as can be seen by comparison with the gray line that is a function proportional to $\sin(\alpha/6)$. b shows the equivalent plot at the energy $E=7.91\,\mathrm{eV}$ where the effect is mostly governed by the sub-band structure leading to pronounced differences of the orbital Edelstein susceptibility between the different nanotubes.
@@
In Fig. 4, we systematically compare all metallic nanotubes with a circumference smaller than $14a\approx 3.4\,\mathrm{nm}$ (radius smaller than $\approx 0.55\,\mathrm{nm}$). In this diagram, each tube is resembled by a hexagon that is placed in a polar coordinate system: the distance from the origin is the tube’s circumference $C$ and the polar angle is the chirality angle $\alpha$. The color of each hexagon resembles the normalized orbital Edelstein susceptibility $\chi_{z}^{L_{z}}\cdot a/T$ at an energy of $0.49\,\mathrm{eV}$ where the considered nanotubes exhibit always only 4 almost linear bands. The magnitude of $\chi_{z}^{L_{z}}\cdot a/T$ takes values of up to ±$\pm 2.2\times 10^{-9}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$ for the considered cases.
@@
We find that $\chi_{z}^{L_{z}}\cdot a/T$ increases with the circumference of the tube $C$ and exhibits an antisymmetric dependence on the chirality angle $\alpha$. For the achiral cases, there is no orbital Edelstein effect: the zigzag tubes $(n,0)$ are characterized by $\alpha=0^{\circ}$ and the armchair tubes $(n,n)$ and $(2n,-n)$ by $\alpha=\pm 30^{\circ}$. The configurations that deviate the most from the two achiral configurations with a preserved inversion symmetry are chiral tubes characterized by $\alpha=\pm 15^{\circ}$. We see that $|\chi_{z}^{L_{z}}|\cdot a/T$ is largest near these angles. In the background of Fig. 4 we have added a contour plot that resembles a function proportional to $C\sin(\alpha/6)$ that approximates the calculated values of $\chi_{z}^{L_{z}}\cdot a/T$ well. The function accounts for the 6-fold rotational symmetry of the graphene lattice.
@@
To better understand the chirality dependence of the orbital Edelstein susceptiblity, we plot this quantity per atom $\chi_{z,\mathrm{atom}}^{L_{z}}=\chi_{z}^{L_{z}}\cdot\frac{1}{2}\cdot a^{2}\frac{\sqrt{3}}{2}/(C\cdot T)$ in Fig. 5a. We can nicely see a $\sin(\alpha/6)$ dependence which indicates that the effect is purely governed by the structural chirality of the nanotube in this energy range. For example, both the $(5,2)$ and $(10,4)$ nanotubes are characterized by the same chirality angle $\alpha\approx-16.10^{\circ}$ and give rise to nearly the same $\chi_{z,\mathrm{atom}}^{L_{z}}$ (indicated by arrows). However, it is important to remember that $\chi_{z}^{L_{z}}\cdot a/T$ is the relevant quantity to compare different nanotubes and this quantity is twice as large for the $(10,4)$ nanotube compared to the $(5,2)$ nanotube due to the double circumference $C$.

### v2

@@ Abstract
CNTs can be constructed mathematically by ‘cutting’ a rectangle out of the two-dimensional carbon allotrope graphene and rolling it to a cylinder. The rotation angle $\alpha$ between the rectangle and the underlying graphene lattice influences the resulting physical properties; most famously, some tubes are metallic and others are insulating. While the two most well-known types of CNTs – tubes with armchair or zigzag edges – posses inversion symmetry, all other nanotubes characterized by $0^{\circ}<|\alpha|<30^{\circ}$ have a broken inversion symmetry which makes them chiral. Over the recent years, such chirality-pure nanotubes have become more easily available (cf. review [4]), such as the chiral $(12,6)$ CNTs that were grown with a Co7W6 or Mo2C catalyst with a purity of more than $90\%$ [13, 14]. Furthermore, dispersion by surfactants [15] and other methods [4] allow for sorting grown tubes based on their chirality.
Figure 1: Chiral carbon nanotube. a Geometrical construction of a $(n,m)=(7,1)$ nanotube from a two-dimensional graphene lattice. The color indicates the $x$ coordinate of each atom. The circumferential vector $\mathbf{C}$ and the translational vector $\mathbf{T}$ span a rectangle containing the atoms of the unit cells of the corresponding chiral carbon nanotube in b. The chirality angle $\alpha$ is indicated. c Top view of the nanotube.
@@
Figure 2: Chirality-induced orbital selectivity of a chiral carbon nanotube. a Band structure of a metallic $(n,m)=(7,1)$ nanotube for which the color indicates the value of the orbital angular momentum $L_{\nu,z}(k)$ normalized to a unit tube length (see legend). b Orbital Edelstein susceptibility $\chi_{z}^{L_{z}}\cdot a/T$ as a function of energy for a $(n,m)=(7,1)$ nanotube (blue) and for the corresponding nanotube with opposite chirality (n,m)=(8,1)(n,m)=(8,-1) (red). The dashed line indicates an energy of $E=0.49\,\mathrm{eV}$ in the range where the orbital Edelstein susceptibility increases quadratically with energy. c Band structure and d orbital Edelstein susceptibility (blue) of an insulating $(n,m)=(4,2)$ nanotube and for the corresponding nanotube with opposite chirality (n,m)=(6,2)(n,m)=(6,-2) (red).
Over the past few years, the orbital Hall effect [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] and the orbital Edelstein effect [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56] have been established as the counterparts to the spin Hall effect [57, 58] and spin Edelstein effect [59, 60]. It has been calculated that the orbital Edelstein effect might explain the large measured CISS: It might mostly be caused by chirality-induced orbital selectivity (CIOS) that is only converted to CISS upon measurement [47, 48, 53, 54, 61, 56, 62]. The orbital angular momentum is transformed to spin angular momentum due to the spin-orbit coupling in the electrodes. In a recent study, we have shown that CIOS in a helix structure like in tellurium has a purely geometrical origin and can be explained even classically based on a helical trajectory of a moving electron [62]. Orbital angular momentum is generated by a rotational motion and is transported along the helix. The effect is caused by intersite-hybridization of the wave function and has to be calculated using the modern formulation [63, 64, 65, 66, 67] instead of relying on the often-used atomic-center approximation. So far, the effect has not been analyzed for more complicated structures, like CNTs, where the sub-band structure and Dirac physics make a classical comparison questionable.
@@
Here we consider ideal CNTs that are periodic along the $z$ direction. The unit cell can be constructed geometrically by rolling up a rectangular cut of a two-dimensional graphene layer. This layer is characterized by two basis atoms and the two lattice vectors $\mathbf{a}_{1}=a\mathbf{e}_{x}$ and $\mathbf{a}_{2}=(a/2)\mathbf{e}_{x}-(a\sqrt{3}/2)\mathbf{e}_{y}$ with the lattice constant a=2.46a=2.46\,Å.
@@
The rectangle that ultimately forms the unit cell of the nanotube is spanned by the circumferential vector $\mathbf{C}=n\mathbf{a}_{1}+m\mathbf{a}_{2}$ and the translational vector $\mathbf{T}=\frac{2m+n}{d}\mathbf{a}_{1}-\frac{2n+m}{d}\mathbf{a}_{2}$ with $d$ the greatest common divisor of $2m+n$ and $2n+m$. The values of $n\in\mathbb{N}^{+}$ and $m=0,1,\dots,n$ determine the configuration of the tube. If the tube was cut horizontally, a $(n,m)=(n,0)$ tube would have a zigzag edge and a $(n,m)=(n,n)$ tube would have an armchair edge. Both of these types of tubes are achiral but all other tubes, for which $m\neq 0$ and $m\neq n$, are chiral. Note that tubes with the opposite chirality can be constructed by allowing also for negative values of $m$.
@@
Figure 3: Band structure of a metallic carbon nanotube by backfolding the graphene band structure. a Brillouin zone of two-dimensional graphene (hexagon) and smaller Brillouin zone (gray) corresponding to the unit cell spanned by $\mathbf{C}$ and $\mathbf{T}$ (gray rectangle in Fig. 1a) for a $(n,m)=(7,1)$ tube. Black lines represent multiples of the reciprocal lattice vector $\mathbf{b}_{T}$ shifted by multiples of the reciprocal lattice vector $\mathbf{b}_{C}$. b Band structure of two-dimensional graphene. The black lines from panel a give rise to the band structure of the corresponding nanotube in c once they are shifted back to the small Brillouin zone (gray). The color resembles $L_{\nu,z}^{\mathrm{approx}}$ that approximates $L_{\nu,z}(k)$ of the corresponding carbon nanotube. The quantity has been normalized by $a/T$ for comparability. Colored dots indicate the Dirac points originally located at the points $K$ and $K^{\prime}$ of the Brillouin zone of graphene.
@@
An example of the geometrical construction of a chiral CNT is shown in Fig. 1. The circumferential vector shown in 1a is $\mathbf{C}=7\frac{1}{2}a\mathbf{e}_{x}-(a\sqrt{3}/2)\mathbf{e}_{y}=7\mathbf{a}_{1}+1\mathbf{a}_{2}$. The gray rectangle is rolled up to form the unit cell of the $(7,1)$ CNT in Fig. 1b. This nanotube is chiral with a chirality angle of
@@
$\alpha=\sphericalangle(\mathbf{T}_{\mathrm{graphene}},\mathbf{e}_{y})=\arctan\frac{-m\sqrt{3}/2}{n+m/2},$ (1)
@@
in our example $\alpha\approx-6.59^{\circ}$. The angle can be recognized by following the color code in Fig. 1b along the translational direction $\mathbf{T}$. When we look into the tube from the top (cf. Fig. 1c), this color code results in a spiral as a signature of the chirality. Due to the 6-fold rotational symmetry of graphene, we will analyze only configurations with $|\alpha|\leq 30^{\circ}$ and focus on the ones with negative chirality angles. Note, that the definition of the chirality angle $\alpha$ varies throughout the literature and can be defined with the opposite sign by choosing a different reference line of the graphene lattice.
@@
Diagonalizing the tight-binding Hamiltonian results in the band structure and eigenvectors. The band structure is symmetric and exhibits many bands $E_{\nu}(k)$ due to the large unit cell. Since the tube is only periodic along the tube vector $\mathbf{T}\parallel\mathbf{e}_{z}$, we find dispersion only along $k\equiv k_{z}$. The system behaves quasi-one-dimensional. Note, however, that the chirality of the system is accounted for by the hopping paths along the $x$ and $y$ directions.
@@
The band structure of our example nanotube $(n,m)=(7,1)$ (cf. Fig. 2a) consists of 76 bands and can be easily understood based on backfolding the graphene band structure [68, 69] ; cf. Fig. 3. Most importantly, the two Dirac points at the Fermi level of graphene exist also for this nanotube. This is the case for all metallic nanotubes, $\mathrm{mod}(n-m,3)=0$. All other nanotubes exhibit a band gap and have no Dirac points near the Fermi level and are characterized by $\mathrm{mod}(n-m,3)=1,2$; an example is shown in Fig. 2c.
@@
In Fig. 2(a,c), the orbital angular momentum is added as a colorcode (normalized to a comparable length, $L_{\nu,z}(k)\cdot a/T$). It has been calculated based on the modern formulation of orbital magnetization (for details see Methods). For the considered metallic nanotube, the values are in the order of up to $\sim\pm 0.5\,\hbar$ and depend antisymmetrically on $k$. $L_{\nu,z}$ is exactly zero at the energy of the Dirac point. In fact, if we expand the system linearly around this energy, the orbital angular momentum is zero everywhere $L_{\nu,z}(k)\equiv 0$. This can be interpreted as a consequence of the massless Dirac fermions. However, the actual band structure deviates from a perfect Dirac cone. As a consequence, $L_{\nu,z}(k)$ increases approximately linearly with $(k-k_{\mathrm{Dirac}})$. However, this dependence is slightly different for $k>k_{\mathrm{Dirac}}$ and $k<k_{\mathrm{Dirac}}$. For the other Dirac point, the orbital angular momentum is exactly reversed. This delicate $k$-dependence determines the orbital Edelstein effect in this energy range, as we will discuss later.
@@
The orbital angular momentum is caused by the circular motion in the nanotube. In equilibrium, it can even be determined analytically based on topological arguments according to Refs. [70, 71]. Coming back to the two-dimensional graphene layer that forms the nanotube, this motion corresponds to a translational motion along the direction $\mathbf{C}$. Making use of the semiclassical formula $\mathbf{L}=\mathbf{r}\times\mathbf{p}$, we find for each band of the nanotube that $L_{\nu,z}$ is approximately given by the product of the radius of the tube $r=\nicefrac{{C}}{{2\pi}}$, the electron mass $m_{e}$ and the velocity $v_{C}=\frac{1}{\hbar}\frac{\partial E_{\mathrm{2d}}(k_{T},k_{C})}{\partial k_{C}}$ along the $\mathbf{C}$ direction which is calculated from the band structure $E_{\mathrm{2d}}(k_{T},k_{C})$ of the graphene layer as $L_{\nu,z}^{\mathrm{approx}}(k)\sim m_{e}rv_{C}(k_{T},k_{C})=\frac{m_{e}}{\hbar}\frac{C}{2\pi}\frac{\partial E_{\mathrm{2d}}(k_{T},k_{C})}{\partial k_{C}}$. Note that $k_{T}$ corresponds to $k$ in the nanotube and $k_{C}$ determines the sub-band index $\nu$.
@@
The orbital Edelstein susceptibility per unit length of the tube $\chi_{z}^{L_{z}}\cdot a/T$ (Fig. 2b) is considerably large in that energy range with values up to $\sim 10^{-8}\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. At $E=0.49\,\mathrm{eV}$, (dashed line in Fig. 2) $\chi_{z}^{L_{z}}\cdot a/T\sim 10^{-9}\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. For energies farther away from the location of the Fermi level $E_{\mathrm{F}}=0$, we find oscillations with energy. Here, the orbital Edelstein susceptibility even reaches values of $\sim 10^{-7}\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$. The oscillations can be understood based on the formation of the sub-band structure by backfolding, shown in Fig. 3. The backfolding leads to different oscillations of the orbital Edelstein susceptibility for different metallic CNTs except for the energy region close to the Fermi level. Here, there are always only four bands that resemble the two Dirac points. The orbital Edelstein susceptibility for different metallic CNTs, including the (12,6) CNTs mentioned in the introduction, is shown in Fig. S1a in the Supplementary Information and will be systematically analyzed below.
Figure 4: Comparison of metallic carbon nanotubes. The diagram shows all metallic nanotubes that are characterized by a circumference $C\leq 14a$ (plotted along the radial direction) and a chirality angle $|\alpha|\leq 30^{\circ}$ (plotted along the polar direction). The color of the hexagon indicates the value of the orbital Edelstein susceptibility at an energy of $E=0.49\,\mathrm{eV}$ (dashed line in Fig. 2). This value has been normalized by the length of the tube $T$, for comparability: $\chi_{z}^{L_{z}}\cdot a/T$ (see legend). The color in the background is the function proportional to $C\sin(6\alpha)$.
Here we want to note that for the insulating CNTs, the backfolded band structure, shown in Fig. S2 in the Supplementary Information, reveals that there are no Dirac points and that there is a band gap at the Fermi level. Still, the origin of the finite orbital Edelstein susceptibility is the same as for the metallic cases: We see a finite orbital angular momentum that is antisymmetric with $k$, as presented in Fig. 2c for a $(n,m)=(4,2)$ tube. Like for the metallic cases, there are oscillations in the energy-dependent orbital Edelstein susceptibility (Fig. 2d) that can be explained based on the backfolding. A comparison of various other insulating CNTs is shown in Fig. S1b in the Supplementary Information.
@@
Figure 5: Chirality dependence of the orbital Edelstein susceptibility per atom. a The blue points are the calculated values $\chi_{z,\mathrm{atom}}^{L_{z}}$ at an energy of $E=0.49\,\mathrm{eV}$ for all metallic nanotubes shown in Fig. 4. Here the band structure exhibits four nearly linear bands and the effect is governed by the chirality of the tube as can be seen by comparison with the gray line that is a function proportional to $\sin(6\alpha)$. b shows the equivalent plot at the energy $E=7.91\,\mathrm{eV}$ where the effect is mostly governed by the sub-band structure leading to pronounced differences of the orbital Edelstein susceptibility between the different nanotubes.
@@
In Fig. 4, we systematically compare all metallic nanotubes with a circumference smaller than $14a\approx 3.4\,\mathrm{nm}$ (radius smaller than $\approx 0.55\,\mathrm{nm}$). In this diagram, each tube is resembled by a hexagon that is placed in a polar coordinate system: the distance from the origin is the tube’s circumference $C$ and the polar angle is the chirality angle $\alpha$. The color of each hexagon resembles the normalized orbital Edelstein susceptibility $\chi_{z}^{L_{z}}\cdot a/T$ at an energy of $0.49\,\mathrm{eV}$ where the considered nanotubes exhibit always only 4 almost linear bands. The magnitude of $\chi_{z}^{L_{z}}\cdot a/T$ takes values of up to $\pm 2.2\times 10^{-9}\,\mu_{\mathrm{B}}\,\mathrm{m}/\mathrm{V}$ for the considered cases.
@@
We find that $\chi_{z}^{L_{z}}\cdot a/T$ increases with the circumference of the tube $C$ and exhibits an antisymmetric dependence on the chirality angle $\alpha$. For the achiral cases, there is no orbital Edelstein effect: the zigzag tubes $(n,0)$ are characterized by $\alpha=0^{\circ}$ and the armchair tubes $(n,n)$ and (2n,n)(2n,-n) by $\alpha=\pm 30^{\circ}$. The configurations that deviate the most from the two achiral configurations with a preserved inversion symmetry are chiral tubes characterized by $\alpha=\pm 15^{\circ}$. We see that $|\chi_{z}^{L_{z}}|\cdot a/T$ is largest near these angles. In the background of Fig. 4 we have added a contour plot that resembles a function proportional to $C\sin(6\alpha)$ that approximates the calculated values of $\chi_{z}^{L_{z}}\cdot a/T$ well. The function accounts for the 6-fold rotational symmetry of the graphene lattice.
@@
To better understand the chirality dependence of the orbital Edelstein susceptiblity, we plot this quantity per atom $\chi_{z,\mathrm{atom}}^{L_{z}}=\chi_{z}^{L_{z}}\cdot\frac{1}{2}\cdot a^{2}\frac{\sqrt{3}}{2}/(C\cdot T)$ in Fig. 5a. We can nicely see a $\sin(6\alpha)$ dependence which indicates that the effect is purely governed by the structural chirality of the nanotube in this energy range. For example, both the $(5,2)$ and $(10,4)$ nanotubes are characterized by the same chirality angle $\alpha\approx-16.10^{\circ}$ and give rise to nearly the same $\chi_{z,\mathrm{atom}}^{L_{z}}$ (indicated by arrows). However, it is important to remember that $\chi_{z}^{L_{z}}\cdot a/T$ is the relevant quantity to compare different nanotubes and this quantity is twice as large for the $(10,4)$ nanotube compared to the $(5,2)$ nanotube due to the double circumference $C$.